### 06 Apr 2026 v3

# Diversity, equity, and inclusion for problems in additive number theory

Abstract.
There is a diversity of problems in additive number theory. Equity requires the consideration of less well known problems, and suggests their inclusion in the additive canon. Of particular interest are problems about the sizes of sumsets of finite sets of integers and problems about the arithmetical structure of intersections of sumsets.
Key words and phrases:
Sumsets in semigroups, restricted sumsets, intersections of sumsets, distribution of of sumset sizes, computational complexity, additive number theory, combinatorial number theory
2000 Mathematics Subject Classification:
11B13, 11B05, 11B75, 11P70, 22D99
This is the text of a talk delivered at the New York Number Theory Seminar at the CUNY Graduate Center on February 19, 2026, and at the International Number Theory Conference (Alladi70) at the University of Florida on March 18, 2026.
1. Canonical additive problems
Here are three of the most important and best known research topics in additive number theory:
  1. (1)
    Partitions and the associated algebraic, analytic, and combinatorial methods,
  2. (2)
    Waring’s problem on sums of $k$th powers and the circle method,
  3. (3)
    Additive prime number theory - Goldbach’s conjecture and sieve methods.
These are the canonical additive problems. (Waring and Goldbach were discussed in my book Additive Number Theory: The Classical Bases [15].) I shall ignore these topics and, instead, focus on other less popular and more “diverse” questions that have not yet been or have not always been included in the additive canon. Fairness, or equity, suggests that less well-known and not obviously “important” topics must always be considered candidates for inclusion, because only time and not today’s “experts” decide what is really important.
2. Sum-product: A problem no one cared about
Paul Erdős  [4], in the book Number Theory Day, Proceedings of the Conference Held at Rockefeller University, New York 1976, wrote the following:
Section 7. A new extremal problem
Let $1\leq a_{1}<\cdots<a_{n}$ be a sequence of integers. Denote by $f_{r}(n)$ the minimal number of distinct integers which are the sum or product of exactly $r$ of the $a$’s.
I conjectured that for every $r$ and $\varepsilon>0$ if $n>n(\varepsilon,r)$, $f_{r}(n)>n^{r-\varepsilon}$.
If true this seems extremely difficult and seems to require new ideas (unless of course an obvious point is being overlooked).
Szemerédi and I observed that it follows from deep results of Freiman that
$\lim_{n\rightarrow\infty}f_{2}(n)/n=\infty$
but even the proof of $f_{2}(n)>n^{1+\varepsilon}$ seems to present great difficulties.
Several years later, in 1983, Erdős and Szemerédi [6] proved that there exists an $\varepsilon>0$ such that
$f_{2}(n)>n^{1+\varepsilon}.$
An important result, but did anyone pay attention? In 1991, Erdös [5] wrote that “our paper with Szemerédi has nearly been forgotten.”
In 1997, Nathanson [18] computed the first explicit value for $\varepsilon$ :
$\varepsilon=\frac{1}{31}\operatorname{\qquad\text{and}\qquad}f_{2}(n)>cn^{32/31}$
where $c=0.00028\ldots$. The explicit exponent $\varepsilon$ was quickly improved.
Ford [7] proved
$\varepsilon=\frac{1}{15}-o(1).$
Elekes [3] proved
$\varepsilon=\frac{1}{4}-o(1)$
Solymosi [30] proved
$\varepsilon=\frac{1}{3}-o(1)$
Bloom [2] has the current best result:
$\varepsilon=\frac{1}{3}+\frac{5}{9813}-o(1)$
The “sum-product” problem has become big business in what is now called “additive combinatorics.”
3. Freiman’s theorem
In the 1976 Number Theory Day conference at Rockefeller University, Erdős wrote,
“Szemerédi and I observed that it follows from deep results of Freiman that ….”
Almost no one had paid any attention to Grigorii Freiman’s remarkable work on inverse problems in additive number theory, even though he had published a book in Russian in 1966 and the American Mathematical Society (Freiman [9]) had published an English translation in 1973. (Of course, in Freiman’s book and papers, the proof of his big theorem was hard to understand.)
Much later, various Hungarians worked on the theorem, and Ruzsa [28] found a beautiful proof in 1994, which was exposed in the Springer Graduate Text Additive Number Theory: Inverse Theorems and the Geometry of Sumsets (Nathanson [16]).
Tim Gowers [10] used this to obtain explicit estimates for Szemerédi’s theorem on $k$-term arithmetic progressions in dense sets of integers, and Freiman’s theorem and its consequences became another big business in additive number theory.
4. A noncanonical problem: Sizes of sumsets
In their 1983 paper on the sum-product conjecture, Erdős and Szemerédi [6] wrote:
Let $2k-1\leq t\leq\frac{k^{2}+k}{2}$. It is easy to see that one can find a sequence of integers $a_{1}<\ldots<a_{k}$ so that there should be exactly $t$ distinct integers in the sequence $a_{i}+a_{j},1\leq i\leq j\leq k$.
The lower bound comes from $k$-term arithmetic progressions. Every integer in the interval of integers $\{0,1,2,\ldots,2k-2\}$ is a sum of two integers in the arithmetic progression $A=\{0,1,2,\ldots,k-1\}$. Moreover, the set of 2-fold sums of a set of $k$ integers has size 2k12k-1 if and only if the set is an arithmetic progression.
The upper bound is combinatorial: If $A=\{a_{1},\ldots,a_{k}\}$ is a set of size $k$ whose $\binom{k}{2}$ sums $a_{i}+a_{j}$ with $a_{i}\neq a_{j}$ and $k$ sums $a_{i}+a_{i}=2a_{i}$ are distinct, then the number of 2-fold sums is
$\binom{k}{2}+k=\binom{k+1}{2}=\frac{k^{2}+k}{2}.$
Such sets are called Sidon sets. For example, the set $\{1,3,3^{2},\ldots,3^{k-1}\}$ is a Sidon set.
For every set $A$ of integers, consider the 2-fold sumset
$2A=\{a+a^{\prime}:a,a^{\prime}\in A\}.$
For all positive integers $k$, we define the sumset size set
$\mathcal{R}(2,k)=\left\{\left|2A\right|:A\subseteq\mathbf{Z}\text{ and }|A|=k\right\}.$
Theorem 1.
For all positive integers $k$,
$\mathcal{R}(2,k)=\left[2k-1,\frac{k^{2}+k}{2}\right].$
This is the Erdős-Szemerédi statement.
Problem 1.
For all $t\in\mathcal{R}(2,k)$, compute (construct) a set $A$ such that $|A|=k$ and $\left|2A\right|=t$. Is there an algorithm?
Theorem 2.
For all $t\in\mathcal{R}(2,k)$, there exists a set $A\subseteq\left[0,2^{k}-1\right]$ such that $|A|=k$ and $\left|2A\right|=t$. The set $A$ can be computed in exponential time.
Problem 2.
If there a subexponential or polynomial bound on the set $A$? Do there exist integers $n$ and $c$ such that, for all $k$ and for all $t\in\mathcal{R}(2,k)$, there is a set $A\subseteq\left[0,ck^{n}\right]$ with $|A|=k$ and $\left|2A\right|=t$?
For every set $A$ of integers, let
$\widehat{2A}=\{a+a^{\prime}:a,a^{\prime}\in A\text{ and }a\neq a^{\prime}\}.$
For all positive integers $k$, we define
$\widehat{\mathcal{R}}(2,k)=\left\{\left|\widehat{2A}\right|:A\subseteq\mathbf{Z}\text{ and }|A|=k\right\}.$
Theorem 3.
For all integers $k\geq 2$,
$\widehat{\mathcal{R}}(2,k)=\left[2k-3,\frac{k^{2}-k}{2}\right].$
Moreover, for all $t\in\widehat{\mathcal{R}}(2,k)$, there exists a set $A\subseteq\left[0,2^{k-2}\right]$ such that $|A|=k$ and $\left|\widehat{2A}\right|=t$.
Problem 3.
If there a subexponential or polynomial bound on the set $A$? Does there exist an integer $n$ and $c>0$ such that, for all $k$ and for all $t\in\widehat{\mathcal{R}}(2,k)$, there is a set $A\subseteq\left[0,ck^{n}\right]$ with $|A|=k$ and $\left|\widehat{2A}\right|=t$?
Problem 4.
“Most” sets $A$ of size $k$ are Sidon sets, that is, $|2A|=\binom{k+1}{2}$.
What is the distribution of sumset sizes $|2A|$?
What are the “popular” sizes, and why are they popular?
Problem 5.
[…]
What are the “popular” sizes, and why are they popular?
5. Range of sumset sizes
The $h$-fold sumset of a set $A$ of integers is the set of all sums of $h$ not necessarily distinct elements of $A$:
$hA=\underbrace{A+\cdots+A}_{\text{$h$ summands}}=\left\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i\in[1,h]\right\}.$
The $h$-fold restricted sumset of $A$ is the set of all sums of $h$ distinct elements of $A$:
[…]
$=\left\{\sum_{s\in S}s:S\subseteq A\text{ and }|S|=h\right\}$
where $|S|$ denotes the size (or cardinality) of the set $S$.
The sumset $hA$ and the restricted sumset $\widehat{hA}$ are finite if $A$ is finite.
For all positive integers $h$ and $k$, we define the sumset size set
$\mathcal{R}(h,k)=\mathcal{R}_{\mathbf{Z}}(h,k)=\left\{\left|hA\right|:A\subseteq\mathbf{Z}\text{ and }|A|=k\right\}.$
There is the elementary estimate:
(1) $\mathcal{R}(h,k)\subseteq\left[h(k-1)+1,\binom{h+k-1}{h}\right$].
The lower bound comes from $k$-term arithmetic progressions and the upper bound from $B_{h}$-sets of size $k$.
In general, we consider the range of sumset sizes of finite subsets of an arbitrary additive abelian group or semigroup $X$:
$\mathcal{R}_{X}(h,k)=\left\{\left|hA\right|:A\subseteq X\text{ and }|A|=k\right\}.$
6. Missing sumset sizes for 3-fold sums of sets with 3 elements
From the elementary estimate (1), we obtain
$\mathcal{R}(3,3)\subseteq\{7,8,9,10\}.$
Extensive computer searches showed
$\{7,9,10\}\subseteq\mathcal{R}(3,3).$
Question: Does there exist a set $A$ with $|A|=3$ and $|3A|=8$?
Answer: No.
Question: Why not?
Theorem 4 (Nathanson [19]).
$\mathcal{R}(3,3)=\{7,9,10\}$
Isaac Rajagopal obtained the following beautiful result for 3-fold sumsets:
Theorem 5 (Rajagopal [27]).
For $k\geq 3$,
$\mathcal{R}(3,k)=\{3k-2\}\cup\left[3k,\binom{k+2}{3}\right$].
Problem 6.
For all $k\in\mathbf{N}$,
$\mathcal{R}(4,k)\subseteq\left[4k-3,\binom{k+3}{4}\right$].
Compute $\mathcal{R}(4,k)$ for all $k\in\mathbf{N}$.
The firsrt missing sumset size result was the following:
Theorem 6 (Nathanson [19]).
For all $h\geq 3$ and $k\geq 3$,
$hk-h+1=\min\left(\mathcal{R}(h,k)\right)$
but
$hk-h+2\notin\mathcal{R}(h,k).$
Thus, $\mathcal{R}(h,k)$ is never an interval of consecutive integers for $h\geq 3$ and $k\geq 3$. This suggests the question:
Problem 7.
What are other “missing” sumset sizes?
Theorem 7 (Vincent Schinina [29]).
For all $h\geq 3$ and $k\geq 3$,
$hk-h+1\in\mathcal{R}(h,k)\operatorname{\quad\text{and}\quad}hk\in\mathcal{R}(h,k)\operatorname{\quad\text{and}\quad}$
but
$hk-h+2,hk-1]\cap\mathcal{R}(h,k)=\emptyset.$
The following result is fundamental.
Theorem 8 (Rajagopal [27]).
Let
m=h(k1)+1m=h(k-1)+1
and
$\Delta_{h,k}=\bigcup_{\ell=1}^{\min(h,k)-3}[m+\ell h+1,m+\ell h+(h-2-\ell)].$
For all $h$ and $k$,
$\mathcal{R}(h,k)\cap\Delta_{h,k}=\emptyset.$
For all $h$, there exists $k_{h}$ such that, for all $k>k_{h}$,
$\mathcal{R}(h,k)=\left[h(k-1)+1,\binom{h+k-1}{h}\right]\setminus\Delta_{h,k}.$
Problem 8.
Compute $\mathcal{R}(4,k)$ for all $k\geq 4$.
7. Computational complexity
Let $h\geq 2$ and $k\geq 2$. Because the sets $\mathcal{R}(h,k)$ and $\widehat{\mathcal{R}}h,k)$ are finite, there exists an integer $N$ such that, if $t\in\mathcal{R}(h,k)$ or $\widehat{\mathcal{R}}(h,k)$, then there exists $A\subseteq[0,N]$ with $|A|=k$ and $|hA|=t$.
Problem 9.
Compute an integer $N$ with this property.
Problem 10.
Compute the smallest integer $N(h,k)$ with this property.
Noah Kravitz and Nathanson obtained an upper bound.
Theorem 9 (Nathanson [20]).
There is the upper bound
$N(h,k)<4(4h)^{k-1}.$
For fixed $k$, this is polynomial in $h$. For example, $N(h,4)<256h^{3}$.
For fixed $h$, this is exponential in $k$. For example, $N(4,k)<4^{2k-1}$.
Problem 11.
Improve these bounds.
Problem 12.
If $A$ is a set of size $k$, then the sumset $hA$ and the size of the sumset $hA$ can be computed in polynomial time, and so the problem of determining the sumset size set $\mathcal{R}(h,k)$ is in the complexity class NP for all $h$ and $k$.
Prove (or disprove): This problem is neither NP-hard nor NP-complete.
8. Sumset races
Here is another class of sumset size problems in additive number theory.
Let $A_{1}$, $A_{2}$, …, $A_{n}$ be $n$ sets of $k$ integers. For every positive integer $h$, consider the $n$-tuple of $h$-fold sumset sizes
$(|hA_{1}|,|hA_{2}|,\ldots,|hA_{n}|).$
If the sumset sizes are pairwise distinct, we can associate to this $n$-tuple a permutation of the integers $1,2,\ldots,n$ which describes the relative order of the sequence of sumset sizes.
If the sumset sizes are not pairwise distinct, we can still associate to this $n$-tuple an $n$-tuple of positive integers that describes the relative order of the sequence of sumset sizes.
Problem 13.
Given $H$ not necessarily distinct $n$-tuples of positive integers $\tau_{1},\tau_{2},\ldots,\tau_{H}$, do there exist finite sets $A_{1},A_{2},\ldots,A_{n}$ such that, for all $h=1,2,\ldots,H$, the relative of the order of the sumset size $n$-tuple
$(|hA_{1}|,|hA_{2}|,\ldots,|hA_{n}|).$
is the same as the relative order of the $h$-th integer $n$-tuple $\tau_{h}$?
One can ask the same question about sumsets in arbitrary additive abelian groups.
By theorems of Nathanson (for integers) and Khovanskii (for groups), one cannot prescribe infinite sequences of $n$-tuples in sumset races.
After preliminary work by Paul Péringuey and Anne de Roton [25], Noah Kravitrz proved the following;
Theorem 10 (Kravitz [12]).
Let $\tau_{1},\tau_{2},\ldots,\tau_{H},\tau_{\infty}$ be $n$-tuples of positive integers. There exist finite sets $A_{1},\ldots,A_{n}$ such that the sumset size $n$-tuple
$(|hA_{1}|,|hA_{2}|,\ldots,|hA_{n}|).$
has the same relative order as the $n$-tuple $\tau_{h}$ for all $h=1,\ldots,H$, and the sumset size $n$-tuple has the same relative order as the $n$-tuple $\tau_{\infty}$ for all $h>H$.
Corollary 1.
If $\varepsilon_{h}=1,0,$ or 1-1 for all $h=1,\ldots,H$, then there exist finite sets $A$ and $B$ such that
|hA||hB||hA|-|hB|
has the same sign as $\varepsilon_{h}$ for all $h=1,\ldots,H$.
There is a continuous analogue of Theorem 10.
Theorem 11 (Nathanson [23]).
Let $G$ be a locally compact abelian group that is not a torsion group and let $\mu$ be a Haar measure on $G$. For integers $n\geq 2$ and $H\geq 2$ and for any $n$-tuples $\mathbf{u}_{1},\ldots,\mathbf{u}_{H}\in\mathbf{N}^{n}$, there exist measurable subsets $A_{1},\ldots,A_{n}$ of $G$ such that
$\tau\left(\mu(hA_{1}),\ldots,\mu(hA_{n})\right)=\tau(\mathbf{u}_{h})$
for all $h=1,\ldots,H$.
There is the following refinement by Jacob Fox, Noah Kravitz, and Shengtong Zhang.
Theorem 12 (Fox, Kravitz, Zhang [8]).
For every sequence $t_{1},\ldots,t_{H}$ of integers, there exist finite sets $A$ and $B$ such that
$|hA|-|hB|=t_{h}$
for all $h=1,\ldots,H$.
This is true not just for finite sets of integers but for finite subsets of infinite or of sufficiently large abelian groups.
There is a continuous analogue for Lebesgue measurable subsets of $\mathbf{R}$.
9. Lebesgue measurable sets
Theorem 13 (Nathanson).
For integers $n\geq 2$ and $H\geq 2$, let $m_{i,h}\in\mathbf{Z}$ for all $i=1,\ldots,n-1$ and $h=1,\ldots,H$. For all $\theta\in\mathbf{R}$, $\theta>0$, there exist Lebesgue-measurable sets of real numbers $A_{1},\ldots,A_{n}$ such that
$\mu(hA_{i})-\mu(hA_{i+1})=\theta m_{i,h}$
for all $i=1,\ldots,n-1$ and $h=1,\ldots,H$.
Note that, for all integers $j,k$ with $1\leq j<k\leq n$, we have
$\mu(hA_{j})-\mu(hA_{k})=\sum_{i=j}^{k-1}\left(\mu(hA_{i})-\mu(hA_{i+1})\right)$
and so the $\binom{n}{2}$ differences $\mu(hA_{j})-\mu(hA_{k})$ are determined by the n1n-1 successive differences $\mu(hA_{i})-\mu(hA_{i+1})$ for $i=1,\ldots,n-1$.
Problem 14.
For integers $n\geq 2$ and $H\geq 2$, let $\theta_{i,h}\in\mathbf{R}$ for all $h=1,2,\ldots,H$ and $i=1,2,\ldots,n-1$. Do there exist Lebesgue-measurable sets $A_{1},\ldots,A_{n}$ of real numbers (or, in a locally compact additive abelian group, Haar-measurable sets $A_{1},\ldots,A_{n}$) such that
$\mu(hA_{i})-\mu(hA_{i+1})=\theta_{i,h}$
for all $h=1,2,\ldots,H$ and $i=1,2,\ldots,n-1$?
By the previous Theorem, the answer is “yes” if there exist integers $m_{i,j}$ and a real number $\theta$ such that $\theta_{i,h}=\theta m_{i,h}$ for all $i$ and $h$. What is the answer in the general case?
10. Sumset intersection limits
Let $\mathbf{N}=\{1,2,3,\ldots\}$ be the set of positive integers. A decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if $A_{q}\neq A_{q+1}$ for infinitely many $q\in\mathbf{N}$. A decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ that is not asymptotically strictly decreasing is eventually constant.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty subsets of an additive abelian semigroup $X$. We define
$A=\lim_{q\rightarrow\infty}A_{q}=\bigcap_{q=1}^{\infty}A_{q}.$
Let $hA$ be the $h$-fold sumset of $A$. For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $hA\subseteq hA_{q}$ and
$hA\subseteq\bigcap_{q=1}^{\infty}hA_{q}.$
Equivalently,
$h\lim_{q\rightarrow\infty}A_{q}\subseteq\lim_{q\rightarrow\infty}hA_{q}$
for every sequence $(A_{q})_{q=1}^{\infty}$.
For $x\in X$, the representation function $r_{A,h}(x)$ counts the number of $h$-tuples $(a_{1},\ldots,a_{h})\in A\times\cdots\times A$ such that $x=a_{1}+\cdots+a_{h}$. The representation function determines the sumset: $hA=\{x\in X:r_{A,h}(x)>0\}$.
In the additive semigroup $\mathbf{N}$, we have $r_{\mathbf{N},h}(x)<\infty$ for all $x\in\mathbf{Z}$ and $h\geq 2$.
In the additive semigroup $\mathbf{Z}$, we have $r_{\mathbf{Z},h}(x)=\infty$ for all $x\in\mathbf{Z}$ and $h\geq 2$.
Theorem 14 (Nathanson [22] ).
Let $h\in\mathbf{N}$. Let $X$ be an additive abelian semigroup such that $r_{X,h}(x)<\infty$ for all $x\in X$. Let $A$ be a subset of $X$ and let $(A_{q})_{q=1}^{\infty}$ be an asymptotically strictly decreasing sequence of subsets of $X$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
and $h\in\mathcal{H}(A_{q})$. If $r_{X,h}(x)<\infty$ for all $x\in X$ and $h\geq 2$, then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in\mathbf{N}$ and $\mathcal{H}(A_{q})=\mathbf{N}$.
Theorem 15 (Nathanson [22] ).
If $A$ is a set of nonnegative integers and if $(A_{q})_{q=1}^{\infty}$ is a strictly decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h=1,2,3,\ldots$, that is,
$\mathcal{H}(A_{q})=\mathbf{N}.$
In general, it is not true that
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for all asymptotically decreasing sequences in an additive abelian semigroup. It is not even true in the integers. Diego Marques and Nathanson proved the following:
Theorem 16 (Marques-Nathanson [13]).
For every $h_{0}\geq 2$ there exist sequences $(A_{q})_{q=1}^{\infty}$ such that $\{1,\ldots,h_{0}-1\}\subseteq\mathcal{H}(A_{q})$ but $h_{0}\notin\mathcal{H}(A_{q})$ and also that there exist sequences $(A_{q})_{q=1}^{\infty}$ such that $\{1,h_{0}\}\subseteq\mathcal{H}(A_{q})$ but $\{2,3,\ldots,h_{0}-1\}\cap\mathcal{H}(A_{q})=\emptyset$.
Problem 15.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. Consider the set
$\mathcal{H}(A_{q})$ $=\{h\in\mathbf{N}:hA=\bigcap_{q=1}^{\infty}hA_{q}\}$
$=\left\{h\in\mathbf{N}:h\lim_{q\rightarrow\infty}A_{q}=\lim_{q\rightarrow\infty}hA_{q}\right\}.$
Describe the sets of positive integers that arise in this way.
Problem 16.
Compute the set
$\mathcal{H}^{*}(0)=\left\{\mathcal{H}(A_{q}):\bigcap_{q=1}^{\infty}A_{q}=\{0\}\right\}.$
11. The moral of the story
The problems in this paper may interest only a minority of number theorists and, perhaps, only a minority of additive number theorists and additive combinatorialists, but, as the history of the sum-product conjecture and Freiman’s theorem suggests, inclusion of a diversity of problems into the additive number theory discussion may be not only equitable but even beneficial to the larger mathematical community.
It suggests one of Aesop’s fables [1]. Aesop lived in ancient Greece between 620 and 564 BCE. Herodotus, in the fifth century BCE, mentioned him in his Histories in Book II, Chapter 134, where he referred to him as “Aesop the maker of fables” and writes that he was a slave of Iadmon, a Samian. Aesop’s fables were transmitted orally and not collected and written down until three centuries after Aesop’s death. Stories have been added continuously and are still being added, so it is not clear which of Aesop’s fables were recited by Aesop.
One of the oldest and most famous of Aesop’s fables is variously known as “The Old Man and His Sons” or “The Bundle of Sticks”. It is numbered 53 in the Perry Index [26].
There was an old man who had many sons. The sons were always fighting with each other, and the father kept trying unsuccessfully to get them to stop. One day he collected a large number of sticks, and, with a rope, tied them together into a thick bundle. He asked his sons to break the bundle. One by one they tried and failed. Then the father cut the rope and gave his sons the individual sticks, which they broke with ease. “Working together, you can succeed,” the father told them.
The moral of the fable is: In unity there is strength.
This is a good way to think about the diversity of problems in additive number theory.